# A Delayed Radio Flare Traces Kinetic Energy Injection in the SMBHB Candidate SDSS J143016.05+230344.4

### 07 Mar 2026 v1

@@ Abstract
We present 4.7–22.2 GHz Very Long Baseline Interferometry (VLBI) monitoring of the candidate pre-coalescence supermassive black hole binary SDSS J143016.05+230344.4 ($z=0.08105$) from 2022 February to 2024 February, together with quasi-simultaneous 0.7–16.5 GHz connected-array spectra. At all epochs, the radio emission is dominated by a single unresolved milliarcsecond core with $T_{\rm B}\gtrsim 10^{7}$ K, confining the variable emission to $\lesssim 0.3$ pc. The spectra require two self-absorbed synchrotron components: a persistent low-frequency component with $\nu_{\rm p,steady}\approx 0.74$ GHz and $S_{\rm p,steady}\approx 1.22$ mJy, and a flare component whose turnover evolves from $(6.35\,{\rm GHz},0.18\,{\rm mJy})$ in 2022 February–May to $(8.61\,{\rm GHz},0.38\,{\rm mJy})$ in 2022 December and then to $(5.83\,{\rm GHz},0.25\,{\rm mJy})$ in 2023 March–April. The 15 GHz flare fraction peaks at $\simeq 80\%$ and matches the near-epoch VLBI recovery fraction, showing that the high-frequency brightening arises from a new compact synchrotron component. A second 15.2 GHz VLBI-core brightening is detected from 2023 September to 2024 February while the source remains unresolved. Equipartition scalings imply characteristic radii of $R_{\rm eq}\sim 5\times 10^{-4}$ pc for the flare and $\sim 9\times 10^{-3}$ pc for the steady component, and a steep inner circumnuclear density profile, $n\propto R^{-1.7}$. The delayed radio peak is consistent with dissipation of an outflow or jet-base disturbance in a structured circumnuclear medium, while a uniform free–free absorber is disfavored.
Active galactic nuclei (16); Accretion (14); Radio jets (1347)
software: CASA (v5.3.0 and v5.6.1; McMullin et al., 2007), AIPS (Greisen, 2003), DiFX software correlator (Deller et al., 2011), DIFMAP (Shepherd et al., 1994)
I Introduction
@@
Supermassive black hole binaries (SMBHBs) on the verge of coalescence are laboratories for studying non-stationary accretion: gas streams, shocks, and magnetic flux are forced to respond on dynamical timescales, while gravitational radiation steadily drains angular momentum. This is also the electromagnetic regime most relevant for low-frequency gravitational-wave astronomy—the population targeted today by Pulsar Timing Arrays and, in the coming decade, by LISA (d’Ascoli et al., 2018; Bowen et al., 2018; Bogdanović et al., 2022; Agazie et al., 2023; Amaro-Seoane et al., 2023). In parallel, James Webb Space Telescope (JWST) is pushing direct observations of rapid black-hole growth and dense nuclear gas reservoirs to high redshift, demonstrating that mergers and violent accretion episodes should be common across cosmic time. The difficulty lies in the fact that the decisive physics takes place on sub-parsec scales, where most diagnostics are either unresolved or heavily absorbed. Radio interferometry is one of the few tools that can directly localize the dissipative region and convert a radio spectrum into size, magnetic field, and energy scales.
@@
II Observations and data reduction
II.1 VLBI observations and data reduction
@@ II.2 VLA and uGMRT observations
For the time-resolved spectrum analysis, we group connected-array measurements into three quasi-/near-epoch bins: Epoch I: 2022 Feb–May (uGMRT March + VLA April–May), Epoch II: 2022 Dec (VLA + uGMRT), Epoch III: 2023 Mar–Apr (VLA). These bins are designed to track months-scale evolution while keeping each spectrum sufficiently sampled. VLBI flux densities are not used in the SSA fitting, because VLBI resolves out larger-scale steep-spectrum emission; instead, VLBI is used to localize the flare and constrain compactness.
III Results
III.1 A single unresolved sub-parsec VLBI core
@@
Visibility-domain modeling (Section II) places tight size limits on this core. Updating our early 1.7 GHz constraints (An et al., 2022c), we obtain FWHM upper limits of <0.19<0.19 mas (<0.29<0.29 pc) at 5 GHz and <0.22<0.22 mas (<0.37<0.37 pc) at 15 GHz. The corresponding lower limits on the brightness temperature span $T_{\rm B}\gtrsim(0.5$$36)\times 10^{7}$ K, well in the non-thermal regime expected for radio-quiet AGN(e.g., Blundell and Beasley, 1998; Anderson et al., 2004; Alhosani et al., 2022; Wang et al., 2023a, b; Cheng et al., 2025) and difficult to reconcile with purely thermal star-formation processes. The relative astrometry is consistent with a stationary centroid (Fig. 4). Over the $\Delta t_{\rm obs}\simeq 1.7$ yr baseline, the absence of a separated compact knot implies an illustrative apparent-speed limit of $\lesssim 0.5$0.8c0.8\,c (Appendix C.5); the long-baseline size evolution yields an even tighter constraint, $\lesssim 0.16\,c$, if interpreted as unresolved expansion.
@@
III.2 High-frequency brightening originates in the VLBI core
@@
III.3 Steady+flare component decomposition
@@
and model both terms as homogeneous synchrotron self-absorption (SSA) spectra (Appendix C.2). The steady component is constrained by the best-sampled early data and held fixed across epochs; the flare component is allowed to vary.
@@
IV Physical Interpretation
The VLBI localization and the time-resolved spectral decomposition together reveal a simple phenomenology: a long-lived, low-frequency synchrotron component persists throughout 2022–2023, while a new, compact, self-absorbed component appears first at high frequency, brightens toward late 2022, and then fades and softens. Below we convert that phenomenology into physical constraints on the emitting region and its environment, and we discuss what kinds of nuclear outflow can satisfy the strict compactness constraints.
IV.1 Physical scales and a steep CNM density profile
@@
Because $k$ depends mainly on ratios, it is comparatively insensitive to uncertain absolute normalizations (geometry, modest Doppler factors, or microphysical parameters that shift both components similarly). The robust inference is therefore geometric: the compact flare evolves inside a rapidly rising CNM density toward $\sim 10^{-4}$$10^{-2}$ pc scales.
@@
IV.2 What does the flare evolution imply?
@@
IV.3 Compact scenarios consistent with the constraints
@@
IV.4 Radio–X-ray/optical connection and implications for a putative pre-coalescence SMBHB
@@
The observed sequence, from a dramatic high-energy disturbance to a compact high-frequency-peaking radio flare a few months later, has close phenomenological analogues in changing-look AGN (Ricci and Trakhtenbrot, 2023; Meyer et al., 2025) and in outflow-powered tidal disruption events (e.g., Alexander et al., 2016; Cendes et al., 2021). J1430+2303 shows that similar radiative–kinetic coupling can occur in a radio-quiet nucleus and remain confined to sub-parsec scales. If future monitoring reveals recurrent flares tied to repeat optical/X-ray activity, their cadence and spectral evolution will offer a purely electromagnetic way to track repeated energy injection in a system that has been discussed as an SMBHB candidate. Regardless of the underlying driver, the combination of VLBI localization and broadband turnover evolution provides a concrete observational template for identifying and interpreting rapidly changing galactic nuclei in the era of low-frequency gravitational-wave astronomy.
V Summary
@@
    The months-long delay relative to the high-energy trigger and the sustained, compact nature of the flare argue against a single, freely expanding impulsive ejection as the sole explanation. The data are instead naturally accommodated by compact scenarios that involve confinement or continued energization: a jet-base plasmon that remains unresolved for years, or an outflow that dissipates through shocks in a dense, structured CNM. A uniform free–free absorber is disfavored by the survival of the steady low-frequency component; if external absorption contributes, it must be patchy and partially covering. Continued multi-frequency VLBI and polarization monitoring (core shift, rotation-measure variability, and the possible appearance of a slow knot) provide direct, falsifiable tests.
@@
Table 1: The observation information.
Code date $\nu$ bandwidth $\tau_{\rm obs}$ Antenna Ref.
Y-M-D (GHz) (MHz) (hour)
RA005 2022-02-27 1.7 256 5.5 EVN (14 telescopes) 1
BA154a 2022-03-02 1.6 512 4.5 SC,FD,LA,BR,HN,NL,PT,KP,OV,MK 1
BA154b 2022-03-02 4.9 512 4.5 SC,FD,LA,BR,HN,NL,PT,KP,OV,MK 1
BA157b 2022-04-26 4.7 512 4.0 BR,FD,HN,KP,LA,MK,NL,OV,PT,SC 2
BA157b 2022-04-26 7.6 512 4.0 BR,FD,HN,KP,LA,MK,NL,OV,SC 2
BA157a 2022-05-08 15.2 1024 6.0 BR,FD,KP,MK,NL,OV,LA 2
BA157a1 2022-05-31 15.2 1024 6.0 BR,FD,KP,MK,NL,OV,PT,SC 2
RA006 2023-02-16 22.2 512 6.0 EVN (10 telescopes) 2
BA166a 2023-09-01 15.2 1024 6.0 BR,FD,HN,KP,LA,MK,NL,OV,PT 2
BA166b 2023-12-02 15.2 1024 6.0 BR,FD,HN,KP,LA,MK,NL,OV,PT,SC 2
BA166c 2024-02-18 15.2 1024 6.0 BR,FD,HN,KP,LA,MK,NL,OV,PT 2
Notes: Col. 1 – Project code; Col. 2 – Observation date; Col. 3 – Central frequency; Col. 4 – The bandwidth used for the imaging; Col. 5 – The observation time; Col. 6 – The telescopes participated in the corresponding observations. The complete VLBA antennas and their name abbreviations are BR (Brewster), FD (Fort Davis), HN (Hancock), KP (Kitt Peak), LA (Los Alamos), MK (Mauna Kea), NL (North Liberty), OV (Owens Valley), PT (Pie Town) and SC (Saint Croix). The EVN telescopes that participated in the observation RS005 are referred to An et al. (2022c), and the EVN telescopes that participated in RA006 are described in Appendix A. References: 1 – An et al. (2022c); 2 – the present paper.
Table 2: Parameters of J1430+2303 obtained from Difmap and MCMC model fitting.
Code $\nu$ Beam rms noise $S_{\rm peak}$ $S_{\rm total}^{\rm difmap}$ $S_{\rm total}^{\rm mcmc}$ $\theta^{\rm difmap}$ $\theta^{\rm mcmc}$ $T_{\mathrm{B}}$ Ref.
[…]
(1) (2) (3) (4) (5) (6) (7) (8) (9) (10) (11)
[…]
Col. 1 – project code; Col. 2 – central frequency; Col. 3 – restoring beam of the image; Col. 4 – rms noise of the image; Col. 5 – peak flux density of the component; Col. 6 and 7 – total flux density of the component derived from Difmap and MCMC, respectively; Col. 8 and 9 – upper limit of the source size estimated from Difmap and MCMC, respectively; Col. 10 – brightness temperature of the VLBI component; Col. 11 – reference of the data.
Table 3: Summary of connected-array radio observations of J1430+2303.
Code date $\nu$ Beam rms noise $S_{\rm peak}$ $S_{\rm total}$
[…]
VLA/21B-375 2022-04-06 8.5 0.011 $\times$ 0.028, 78 13 326 352 $\pm$ 6
9.5 0.074 $\times$ 0.027, 78 12 316 332 $\pm$ 2
2022-04-13 5.0 0.57 $\times$ 0.32, 69 17 397 418 $\pm$ 3
6.0 0.48 $\times$ 0.26, 70 18 403 412 $\pm$ 10
2022-05-04 2.5 0.73 $\times$ 0.61, -65 21 457 473 $\pm$ 7
3.5 0.55 $\times$ 0.47, -66 16 458 496 $\pm$ 9
2022-05-06 12.7 0.24 $\times$ 0.12, 71 17 258 283 $\pm$ 7
14.3 0.21 $\times$ 0.12, 68 15 226 219 $\pm$ 13
15.7 0.19 $\times$ 0.11, 68 16 205 226 $\pm$ 10
17.3 0.18 $\times$ 0.10, 70 18 160 150 $\pm$ 6
VLA/22B-106 2022-12-26 4.5 4.03 $\times$ 3.69, -3 35 501 522 $\pm$ 6
6.5 2.89 $\times$ 2.59, 7 30 478 476 $\pm$ 8
7.5 2.55 $\times$ 2.27, 12 24 463 486 $\pm$ 13
2.5 2.58 $\times$ 1.86, 161 69 640 720 $\pm$ 10
3.5 5.07 $\times$ 4.61, 4.89 46 590 590 $\pm$ 10
2022-12-27 12.8 0.69 $\times$ 0.43, 111 21 394 432 $\pm$ 10
14.3 1.96 $\times$ 1.26, -63 22 431 435 $\pm$ 8
15.9 1.80 $\times$ 1.15, -65 22 421 416 $\pm$ 16
17.4 1.69 $\times$ 1.10, -61 26 433 461 $\pm$ 23
2023-03-07 15.1 0.15 $\times$ 0.08, 143 12 296 322 $\pm$ 7
10.0 0.55 $\times$ 0.49, 5 14 398 410 $\pm$ 5
2023-03-09 6.0 0.35 $\times$ 0.19, 131 19 449 481 $\pm$ 11
2023-04-26 13.5 0.45 $\times$ 0.40, -56 16 212 226 $\pm$ 7
16.6 0.38 $\times$ 0.33, -64 19 161 202 $\pm$ 19
2023-04-26 9.0 0.62 $\times$ 0.57, -6 18 389 421 $\pm$ 6
11.0 0.52 $\times$ 0.47, 8 19 331 355 $\pm$ 6
5.0 0.40 $\times$ 0.24, 30 22 482 532 $\pm$ 7
7.0 0.29 $\times$ 0.17, 147 21 388 426 $\pm$ 13
GMRT/ddtC244 2022-03-09 1.3 1.30 $\times$ 0.91, 106 37 1002 1185 $\pm$ 23
2022-03-15 0.75 1.49 $\times$ 0.70, 159 46 1115 1292 $\pm$ 45
2022-12-26 1.3 2.04 $\times$ 1.92, 139 23 838 880 $\pm$ 12
2022-12-27 0.75 2.05 $\times$ 1.24, 88 53 1026 1198 $\pm$ 49
VLASS 2017-11-17 3.0 2.51 $\times$ 2.20, 52 116 430 430 $\pm$ 116
2020-07-16 3.0 2.38 $\times$ 2.10, -11 130 441 441 $\pm$ 130
2023-01-23 3.0 2.51 $\times$ 2.20, 52 115 581 581 $\pm$ 115
ASKAP/RACS 2022-01-05 1.7 16.90 $\times$ 8.90, 152 174 683 683 $\pm$ 174
2022-07-29 1.4 21.09 $\times$ 10.60, 2 181 905 905 $\pm$ 181
2020-10-16 0.888 25.00 $\times$ 25.00 299 1031 1031 $\pm$ 299
2020-10-17 0.888 25.00 $\times$ 25.00 299 1044 1044 $\pm$ 299
Table 4: SSA decomposition parameters and derived physical parameters.
Component $\nu_{p}$ (GHz) $F_{\nu,p}$ (mJy) $R_{\rm eq}$ (pc) $B_{\rm eq}$ (G)
Steady $0.740^{+0.090}_{-0.110}$ $1.221^{+0.120}_{-0.080}$ $9.14\times 10^{-3}$ 0.102
Flare (Epoch I) $6.346^{+0.593}_{-0.570}$ $0.176^{+0.024}_{-0.022}$ $4.24\times 10^{-4}$ 1.081
Flare (Epoch II) $8.608^{+0.851}_{-0.744}$ $0.384^{+0.027}_{-0.028}$ $4.53\times 10^{-4}$ 1.351
Flare (Epoch III) $5.831^{+0.646}_{-0.645}$ $0.253^{+0.030}_{-0.025}$ $5.49\times 10^{-4}$ 0.956

### 19 Mar 2026 v2

@@ Abstract
Context. SDSS J143016.05+230344.4 ($z=0.08105$) has been proposed as a candidate pre-coalescence supermassive black hole binary and shows remarkable multiwavelength variability. Its radio evolution provides a direct probe of the compact emitting region and of the physical origin of the late-time activity.
Aims. We aim to localize the variable radio emission, characterize its spectral evolution, and constrain whether the radio brightening is produced by a newly emerging compact component, external absorption, or dissipation in a structured circumnuclear environment.
Methods. We analyzed 4.7–22.2 GHz Very Long Baseline Interferometry (VLBI) observations obtained between 2022 February and 2024 February, together with quasi-simultaneous connected-array spectra covering 0.7–16.5 GHz. We combined the VLBI brightness-temperature and compactness constraints with spectral decomposition and equipartition-based estimates of source size and ambient density structure.
@@
Results. At all epochs, the radio emission is dominated by a single unresolved milliarcsecond core with $T_{\rm B}\gtrsim 10^{7}$ K, constraining the variable emission to $\lesssim 0.3$ pc. The broadband spectra require two synchrotron self-absorbed components: a persistent low-frequency component with $\nu_{\rm p,steady}\approx 0.74$ GHz and $S_{\rm p,steady}\approx 1.22$ mJy, and a flare component whose turnover evolves from $(6.35\,{\rm GHz},0.18\,{\rm mJy})$ in 2022 February–May to $(8.61\,{\rm GHz},0.38\,{\rm mJy})$ in 2022 December, and then to $(5.83\,{\rm GHz},0.25\,{\rm mJy})$ in 2023 March–April. The flare contribution at 15 GHz reaches $\sim 80\%$ and matches the near-epoch VLBI recovery fraction, showing that the high-frequency brightening arises from a newly formed compact synchrotron component. A second brightening of the 15.2 GHz VLBI core is detected between 2023 September and 2024 February, while the source remains unresolved. Equipartition scalings imply characteristic radii of $\sim 5\times 10^{-4}$ pc for the flare and $\sim 9\times 10^{-3}$ pc for the steady component, and indicate a steep inner circumnuclear density profile, $n\propto R^{-1.7}$.
Conclusions. The delayed radio flare is best explained by dissipation in an outflow or jet-base disturbance propagating through a structured circumnuclear medium.
Key Words.:
Galaxies: actives – Galaxies: jets – Galaxies: individual: J1430+2303 – Galaxies – supermassive black holes – Techniques: high angular resolution – Instrumentation: interferometers
1 Introduction
@@
Supermassive black hole binaries (SMBHBs) on the verge of coalescence are laboratories for studying non-stationary accretion: gas streams, shocks, and magnetic flux are forced to respond on dynamical timescales, while gravitational radiation steadily drains angular momentum. This is also the electromagnetic regime most relevant for low-frequency gravitational-wave astronomy—the population targeted today by Pulsar Timing Arrays and, in the coming decade, by LISA (d’Ascoli et al., 2018; Bowen et al., 2018; Bogdanović et al., 2022; Agazie et al., 2023; Amaro-Seoane et al., 2023). In parallel, James Webb Space Telescope (JWST) is pushing direct observations of rapid black-hole growth and dense nuclear gas reservoirs to high redshift, demonstrating that mergers and violent accretion episodes should be common across cosmic time. Characterizing the underlying physics of these systems remains inherently difficult, as the decisive physics takes place on sub-parsec scales, where most diagnostics are either unresolved or heavily absorbed. Radio interferometry is one of the few tools that can directly localize the dissipative region and convert a radio spectrum into size, magnetic field, and energy scales.
@@
2 Observations and data reduction
2.1 VLBI observations and data reduction
@@ 2.2 VLA and uGMRT observations
For the time-resolved spectrum, we group connected-array measurements into three quasi-/near-epoch bins: Epoch I: 2022 Feb–May (uGMRT March + VLA April–May), Epoch II: 2022 Dec (VLA + uGMRT), Epoch III: 2023 Mar–Apr (VLA). These bins are designed to track months-scale evolution while keeping each spectrum sufficiently sampled. VLBI flux densities are not used in the synchrotron self- absorption (SSA) fitting, because VLBI resolves out larger-scale steep-spectrum emission; instead, VLBI is used to localize the flare and constrain compactness.
3 Results
3.1 A single unresolved sub-parsec VLBI core
@@
Visibility-domain modeling (Section 2) places tight size limits on this core. Updating our early 1.7 GHz constraints (An et al., 2022c), we obtain FWHM upper limits of <0.19<0.19 mas (<0.29<0.29 pc) at 5 GHz and <0.22<0.22 mas (<0.37<0.37 pc) at 15 GHz. The corresponding lower limits on the brightness temperature span $T_{\rm B}\gtrsim(0.5$$36)\times 10^{7}$ K, well in the non-thermal regime expected for radio-quiet AGN(e.g., Blundell and Beasley, 1998; Anderson et al., 2004; Alhosani et al., 2022; Wang et al., 2023a, b; Cheng et al., 2025) and difficult to reconcile with purely thermal star-formation processes. The relative astrometry is consistent with a stationary centroid (Fig. 4). Over the $\Delta t_{\rm obs}\simeq 1.7$ yr baseline, the absence of a separated compact knot implies an illustrative apparent-speed limit of $\lesssim 0.5$0.8c0.8\,c (Appendix C.5); the long-baseline size evolution yields an even tighter constraint, $\lesssim 0.16\,c$, if interpreted as unresolved expansion.
@@
3.2 High-frequency brightening originates in the VLBI core
@@
3.3 Steady+flare component decomposition
@@
and model both terms as homogeneous SSA spectra (Appendix C.2). The steady component is constrained by the best-sampled early data and held fixed across epochs; the flare component is allowed to vary.
@@
4 Physical Interpretation
The VLBI localization and the time-resolved spectral decomposition together reveal a simple phenomenology: a long-lived, low-frequency synchrotron component persists throughout 2022–2023, while a new, compact, self-absorbed component appears first at high frequency, brightens toward late 2022, and then fades and softens. In the framework of time-domain radio loudness, this phenomenology is naturally interpreted as the superposition of a prompt compact component and a longer-lived radio-memory component (An 2026, under review). Below we convert that phenomenology into physical constraints on the emitting region and its environment, and we discuss what kinds of nuclear outflow can satisfy the strict compactness constraints.
4.1 Physical scales and a steep CNM density profile
@@
Because $k$ depends mainly on ratios, it is comparatively insensitive to uncertain absolute normalizations (geometry, modest Doppler factors, or microphysical parameters that shift both components similarly). Recent studies of radio-detected TDEs have likewise shown that radio evolution can provide direct constraints on sub-parsec circumnuclear density profiles, with sample-average behavior broadly consistent with Bondi-like scaling and closure-relation analyses yielding similarly steep CNM gradients in many systems (Goodwin and Mummery, 2026; Zhou et al., 2026). The robust inference is therefore geometric: the compact flare evolves inside a rapidly rising CNM density toward $\sim 10^{-4}$$10^{-2}$ pc scales.
@@
4.2 What does the flare evolution imply?
@@
4.3 Compact scenarios consistent with the constraints
@@
4.4 Radio–X-ray/optical connection and implications for a putative pre-coalescence SMBHB
@@
The observed sequence, from a dramatic high-energy disturbance to a compact high-frequency-peaking radio flare a few months later, has close phenomenological analogues in changing-look AGN (Ricci and Trakhtenbrot, 2023; Meyer et al., 2025) and in outflow-powered tidal disruption events (e.g., Alexander et al., 2016; Cendes et al., 2021). J1430+2303 shows that similar radiative–kinetic coupling can occur in a radio-quiet nucleus and remain confined to sub-parsec scales. In this broader sense, J1430+2303 may represent a nominally radio-quiet nucleus caught during a transient radio-bright phase (An 2026, under review). If future monitoring reveals recurrent flares tied to repeat optical/X-ray activity, their cadence and spectral evolution will offer a purely electromagnetic way to track repeated energy injection in a system that has been discussed as an SMBHB candidate. Regardless of the underlying driver, the combination of VLBI localization and broadband turnover evolution provides a concrete observational template for identifying and interpreting rapidly changing galactic nuclei in the era of low-frequency gravitational-wave astronomy.
5 Summary
@@
    The months-long delay relative to the high-energy trigger and the sustained, compact nature of the flare argue against a single, freely expanding impulsive ejection as the sole explanation. The data are instead naturally accommodated by compact scenarios that involve confinement or continued energization: a jet-base plasmon that remains unresolved for years, or an outflow that dissipates through shocks in a dense, structured CNM. A uniform free–free absorber is disfavored by the survival of the steady low-frequency component; if external absorption contributes, it must be patchy and partially covering. Continued multi-frequency VLBI and polarization monitoring (core shift, rotation-measure variability, and the possible appearance of a slow knot) provide direct tests.
Acknowledgements.